OpenStudy (anonymous):

what is the value of 6.9x10 to the second power?

OpenStudy (anonymous):

use a ^ for powers

OpenStudy (anonymous):

10^2 = 100 6.9*10² = 6.9*100 = 690

OpenStudy (anonymous):

so write 6.9x^10

OpenStudy (anonymous):

okay. I've got it. :) Thank you both.